a. Determine the pH during the titration of 31.3 mL of 0.264 M trimethylamine ((CH3)3N, Kb = 6.3×10-5) by 0.264 M HI at the following points. (Assume the titration is done at 25 °C.)
Note that state symbols are not shown for species in this problem.
(a) Before the addition of any HI ____

(b) After the addition of 12.6 mL of HI ____

(c) At the titration midpoint ____

(d) At the equivalence point ____

(e) After adding 43.8 mL of HI ____

b.

The Ksp for BaSO4 is 1.1×10-10 at 25 °C. Calculate the solubility of barium sulfate in pure water in (a) moles per liter and (b) grams per liter.

_____ moles/liter

_____ g/liter
Answer #1

Answer: a) Before addition of any HI, the following reaction can be written as B + H2O --> BH+ + OH-

B = (CH3)3N. The value of kb for above reaction =  6.3×10-5 (given)

From ICE table;

B ---> BH+ + OH-

Init 0.264M 0 0

change -x x x

Equil. (0.264-x) x x

kb = [BH+][OH-]/[B]

= x2/(0.264-x) or x2 + kb.x - 0.264kb = 0 on solving quadratic equation we will get x = 0.004054

Now p(x) ~ pOH = -log(0.004054) =2.39

pH = 14 - pOH = 11.61

b. After the addition of 12.6 mL of HI

total volume = 43.9 mL (acid + base)

number of moles of acid present in 12.6 mL = 0.264/1000 x 12.6 mL = 3.3264 x 10-3 moles (x1)

number of moles of base present in 31.3 mL = 8.2632 x 10-3 moles (x2)

Following the above ICE table (OH- will be used to neutralize HI);

B + H2O   ----> BH+ + OH-

Ini x2     0 0

chan. -x1     x1 x1

equi. x2-x1 0 x1   x1

Using henderson equation,

x2-x1 = 4.9368 x 10-3 moles or [B] = x2-x1/0.0439 L = 0.1124 M

similarly [BH+] = x1/0.0439 = 0.07577 M

pOH = pKb + log [BH+]/[B] = 4.20 + log(0.6741) = 4.02 M

pH = 14.0 - pOH = 9.98

(c) At the titration midpoint ____
At midpoint [B] = [BH+] , Hence using the method involved in part b

pOH = pKb = 4.2 as log term will be zero

pH = 9.8

(d) At the equivalence point ____

[B] = [HI] ; total volume = 31.3 (from B) + 31.3 (from HI) = 62.6 mL

The concentration of BH+ will determine the pH of solution.

So, the concentration of [BH+]  8.2632 x 10-3 (moles of B in 31.3 mL of solution) / (0.0626 L) = 0.132 M

kw = ka.kb from kb for B, we can calculate the ka for BH+ ; where kw = 10-14

ka = 10-14/6.3×10-5 = 1.587 x10-10  

Now the equilibrium reaction would be;

BH+ + H2O --> B + H3O+

ka = x.x/(0.132 - x) = x2 +kb.x - 0.132.kb = 0

x = 4.5759 x 10-6 ~ [H3O+]

pH = -log[H3O+] = 5.3395

e. After adding 43.8 mL of HI ____

The number of moles of HI present = 43.8 - 31.3 mL of acid (remaining part is consumed)

= 12.5 mL   

Moles of HI = 0.264/1000 x 12.5 = 3.3 x 10-3 moles

Total volume = 31.3 + 43.8 = 75.1 mL

[HI] = 3.3 x 10-3 /0.0751 L = 0.0439 M

For acid such as HI, assuming the dissociation to be ~100%

[H+] = 0.0439 M

pH = 1.35

Second Part The Ksp for BaSO4 is 1.1×10-10 at 25 °C

The Ksp relation for BaSO4 = [Ba2+][SO42-] = x.x = x2.

x = sqrt(Ksp) = 1.0488 x 10-5 M (moles/L)

Molar mass of BaSO4 = 233.38 g/mol

Hence, strength of solution =  molarity x molar mass=  2.4476 x 10-3 g/L

Please let me know if you have any doubt by commenting below the answer.

Thanks